how do you solve fore y?
lny+ ln3=x-1

Answer by oscargut(2103)

lny+ ln3=x-1
ln(3y) = x-1
3y = e^(x-1)
y = (e^(x-1))/3
I can help you with more questions at: mthman@gmail.com
Thanks